

********************************************************************
UARS WEEKLY STATUS REPORT - 09 Jan 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS detectors were active from December 13 through December 16
and December 24 through December 27. Energy deposition due to electrons
was determined from the measured X-ray spectrum. The PEM zenith spectrometers
were powered January 4. HEPS-2 microprocessor errors did not clear and the
HEPS-2 unit was shut down. The HEPS-2 unit was restarted on January 5
successfully. This type of problem has been seen before and appears to be a
symptom of being cold for a long period of time. It is not anticipated that
this problem will occur under solar array control. During the time the PEM
zenith spectrometers are powered, PEM produces energy deposition profiles from
electron and proton precipitation.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Happy New Year!
HRDI is currently off. Next scheduled turn-on is Jan. 23.
***********
WINDII
***********
The WINDII team are preparing for the initiation of the "final"
production run, with Version 5 of the production processing
software.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 355 CD-ROM requests as of
09-January-98. There were no new orders received or shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) Prod (1350R-750C=600L)
The CDHF continues to monitor UKMO data. Data for UARS Days 2263 and
2264 has been requested but not yet received.
The CDHF will load November 1997 current day DAAC data in January 1998.
Pending software upgrades: IDL 5.0, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 December 1997 (Orbit 34387, GMT Day 361) through 02
January 1998 (Orbit 34491, GMT Day 002).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 362/2125 34415 AXIS off
WINDII 362/2125 34415 Off
MLS 362/2235 34416 Off
HRDI 363/0010 34417 On
HALOE 363/0200 34418 On
The beta angle reached a maximum angle of 34.0 degrees on 01 January 1998 (DOY
001). The beta angle is now decreasing toward a minimum angle of 0.0 degrees
on 16 January 1998 (DOY 016).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.1 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 degrees C. Additional battery performance data for each battery on
27 December 1997 and 02 January 1998 are provided in Appendix D.
The clock error ranged from -10.2 to +11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 04
% HALOE 01
% HRDI 01
% MLS 02
% PEM 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT 07
TRANSPONDER FREQUENCIES:
Transponder A: 97/343 - 2287.497007
Transponder B: 97/344 - 2287.499121
GROUND SYSTEM OPERATION
Seven (7) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 98/001 show the observatory data loss to be 20 hours, 41
minutes, 24.692 seconds (an increase of 11.776 seconds since last reported).
This is a 0.0376 percent data loss which equals a 99.9624 percent data capture
for the mission.
The increase of 0.544 seconds occurred on Day 353 and resulted from a hit in
the data.
The increase of 0.544 seconds occurred on Day 354 and resulted from a hit in
the data.
The increase of 7.968 seconds occurred on Day 355 and resulted from
insufficient tape recorder overlap.
The increase of 2.720 seconds occurred on Day 365 and resulted from
insufficient tape recorder overlap and a hit in the data.
The FOT is still working out the details of participating in the IP return-link
telemetry test that the network is conducting.
No testing of the new CMS software release (R7B1V2) occurred this week.
Operational processing of level-0 telemetry analysis files on the UMPGAL
computer (new MPG Alpha machine) continues in parallel with the CDHF computer
system. Quicklook/level-0 file transfers between CDHF and UMPGAL failed
completely on 02 January 1998 at approximately 20:00 UTC. This was the only
file transfer problem identified this week.
OTHER SIGNIFICANT EVENTS
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current status
of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 16 January 1998 at
016/13:25:18 UTC.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust will
probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
no change
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 January 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34391 361/0656 97-640 OBC Clk Adj fr 20.6 to 0.00 msec/day
34399 361/1940 97-641 MLS Disabled Comp Htr 18 B side
34415 362/2125 97-642 PEM AXIS Power OFF
34415 362/2125 97-643 WINDII Power OFF
34416 362/2235 97-644 MLS Power OFF
34417 363/0010 97-645 HRDI Power ON
34418 363/0200 97-646 HALOE Power ON
34426 363/1402 97-647 OBC Clk Adj fr 0.00 to 20.6 msec/day
34456 365/1411 97-648 OBC Clk Adj fr 20.6 to 0.00 msec/day
34485 002/1317 98-001 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 January 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
MLSCANACTTMP Y-LO MLS Scan Act. Temperature 34392-3
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
(none)
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 January 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34391 361/07:00 (27 Dec 97) from 20.6 to 00.0 msec/day
34426 363/14:06 (29 Dec 97) from 00.0 to 20.6 msec/day
34456 365/14:17 (31 Dec 97) from 20.6 to 00.0 msec/day
34485 002/13:21 (02 Jan 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 January 1998
27 Dec 1997 - GMT Day 361 Beta = 28.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.73/-3.68 +0.0/ +0.0 18.7 0.000 0.0
2 +44.8/ -5.6 +7.09/+5.47 +32.0/-16.0 24.8 1.023 23.6
3 +44.8/ -11.2 +5.79/+4.51 +31.2/-17.6 24.8 1.024 22.8
02 Jan 1998 - GMT Day 002 Beta = 33.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.6 0.000 0.0
2 +16.8/ -44.8 +6.11/+4.83 +31.6/-16.0 25.1 1.025 22.4
3 +39.2/ -22.4 +4.83/+3.87 +30.0/-16.0 25.1 1.025 21.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 January 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-124 FOT 34426 363/1402 14170 Generic Late Acquisition
97-124 FOT 34433 364/0205 14170 Generic Late Acquisition
97-124 FOT 34438 364/0903 14170 Generic Late Acquisition
97-128 FOT 34442 364/1628 TAC Crashed
97-129 FOT 34444 364/2003 TAC Crashed
97-130 FOT 34457 365/1610 19646 TAC Crashed
98-002 FOT 34465 001/0714 TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- ---------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- ---------------
None this report period